Suppose we wish to lend $70,000,000 for 91 days at LIBOR beginning next September , and that the quoted Eurodollar futures price (based on a 360-day year)is 91.46. How much will we have after invested for 91 days?
Answer #1

Interest rate = 1- ( 91.46 /100) = 8.54%

Interest for 91 days = 8.54% * (91/360) = 2.1587222%

Money that will be repaid after 91 days = 70,000,000 * (1 +2.1587222%) = 71,511,105.56

we have after invested for 91 days = $71,511,105.56

Answer may be slightly different due to rounding off differences.

Any doubt please comment.

If there are options please provide.